14. Blue Bull Inc, has a target debt-equity ratio of 55. Its WACC is 8.1 percent, and the tax rate is 35 percent Required: (a) If the companys cost of equity is 11 percent, what is its pretax cost of debt? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g. 32.16)) Pretax cost of debt (b)lf the atertax cost of debt Is 3.8 percent, what is the cost of equity? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16) Cost of equy OType here to
Answer #1

WACC = rD (1- Tc )*( D / V )+ rE *( E / V )

Where...

rD = The required return of the firm's Debt financing
(1-Tc) = The Tax adjustment for interest expense
(D/V) = (Debt/Total Value)
rE= the firm's cost of equity
(E/V) = (Equity/Total Value)

0.081 = 0.55/(1+0.55) * (1-0.35) * rD + 1/(1+0.55)*0.11

rD = 0.04350 = 4.35%

b.

WACC = 0.081 = 0.55/(1+0.55) * 0.038 + 1/(1+0.55)*rE

rE = 0.1046 = 10.46%